1) A store drops the price they charge for socks by 10% and in response they see sales increase by 10%. What's the price elasticity of socks?

Group of answer choices

a) 1

b) 2

c) 0.5

d) 0

----------

2) When we say we've reached "consumer equilibrium" we mean that we're getting the:

Group of answer choices

a) maximum total utility possible from our budget

b) best marginal utility from our budget

c) same total utility from each product

d) same number of each product

----------------

3) What do we mean by "marginal product" of labor?

Group of answer choices

a) The increases in output resulting from adding one more unit of labor

b)The last unit of output produced by labor at the end of the period

c)Total output divided by total number of units of labor

d) Smallest unit of output produced by total labor units used

--------

4) When examining how much labor to use in the production of a good or service, in which case has "average product" been maximized?

Group of answer choices

a) MP=AP

b) MP

c) MP>AP

d) MP=0

e) MP=Max

--------

5) Categorize each of these pottery business costs, keeping in mind our definitions of explicit and implicit costs. Determine which one is explicit and which one is explicit.

a) Buying raw clay

b) Interest no longer earned on money used to open business  

c) Purchasing a kiln to fire mugs in

d) Renting a storefront (you don't own)

e) Declining a job offer
Answer #1

1. Price elasticity of socks= % change in quantity demanded of socks/% change in price of socks

= 10/10

= 1

Hence, the correct answer is option a.

2. At consumer equilibrium, the consumer's total utility gets maximized given his budget. In algebraic form, consumer equilibirum is

P1/P2= MU1/MU2 (utility-maximization)

Hence, the correct answer is option a.

3. Marginal product of labor is the change in total output by employing one more unit of labor.

Hence, the correct answer is option a.

4. Both AP and MP curves shows similar trend- both curves first rise, reach their maximum and then fall. At the point where AP reached maximum, AP=MP

Hence, the correct answer is option a.